Session: Nov 2001 (session 3)
Wavelength: 5.0 cm


To give feedback on this experiment, click here

Experiment Feedback

Effelsberg (A. Kraus at Tue Nov 13 08:21:50 MET 2001) Minor failures
Data between UT1337-1349 were lost (antenna was slewing).

\(Other\) (andrea at Tue Nov 13 11:21:34 MET 2001) Minor failures

Medicina (A.Maccaferri at Tue Nov 13 11:22:22 MET 2001) Success

Jodrell_Bank (Alastair Gunn at Wed Nov 14 10:25:30 MET 2001) Minor failures
Serious problem with antenna control occurred at about 21:15 UT. Data may be OK up to 21:30 but after that telescope was refusing all source changes. Cause of problem still not clear.

Cambridge (Alastair Gunn at Wed Nov 14 10:26:04 MET 2001) Minor failures
Problems reading procedure library at start of experiment. Problem solved and data good after 13:48 UT. No other known problems.

Onsala (John Conway at Thu Nov 15 10:48:19 MET 2001) Success
No known problems